OpenStudy (anonymous):

Hi i am studying for a precalc test tomorrow and was wondering if someone could help me with inverse functions!

OpenStudy (saifoo.khan):

Where are you facing problems in finding inverses?

OpenStudy (anonymous):

ok lemme give you the equation...

OpenStudy (saifoo.khan):

Okay, so what you basically do is, write it this way: y = 2x /(3x-1) Now you goal is to bring all "x"s on one side of the equation. Can you do it?

OpenStudy (anonymous):

um ok so you would do it 2x/(3x-1) >> 2x/(3x-1) multiplied with (3x-1) for both top and bottom? right??

OpenStudy (anonymous):

and then on the other side you would multiply (3x-1) by y?

OpenStudy (anonymous):

i am just really confused about this whole thing and i feel like im doing this completely wrong:x

OpenStudy (saifoo.khan):

Yes. you're right. You should get something like: y(3x-1) = 2x

OpenStudy (saifoo.khan):

Now solve further.

OpenStudy (anonymous):

ok good! then it would be .... 3xy-x= 2x?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

but then how do u get it simplified from there?

OpenStudy (saifoo.khan):

No.

OpenStudy (saifoo.khan):

3xy -y = 2x You did a little mistake.

OpenStudy (anonymous):

oh ok! how do u get it simplified from there? so that the x's are all on one side though?

OpenStudy (saifoo.khan):

YES! You are a boss at this.

OpenStudy (anonymous):

wait so u dont have to simplify it?

OpenStudy (saifoo.khan):

You have to make x the subject.

OpenStudy (anonymous):

3xy-y/2= x ?

OpenStudy (saifoo.khan):

3xy -y = 2x 3xy - 2x = y Take x common: x(3y - 2) = y FInally: x = y/(3y-2) Get it?P

OpenStudy (anonymous):

oh i see! thank you so much!